# Unifying uncertainties for rotor-like quantum systems

### 04 Mar 2024 v1

@@ Abstract
Quantum rotor represents the second simplest model after the harmonic oscillator with profound interdisciplinary but yet unexplored implications. It overreaches its mechanical meaning with significant consequences and promising applications in, e.g., singular optics, super-conductive circuits with Josephson junction or optimal pulse shaping in time frequency domains at ultimate quantum limit. Quantification of complementarity between angular momentum and angular variable is essential for exploitation of this canonical pair in quantum metrology. Whereas the natural measures for position and momentum are variances, the uncertainty associated with angular variable can be linked to moments of inertia of the unit ring about axes passing through its center of mass. This interpretation provides variants for saturable uncertainty relations which can be further used in quantum metrology of the quantum rotor and explains ambiguities in choice of possible uncertainty measures associated with sine and cosine functions of angular variable. Special attention will be payed to uncertainty products which are exactly or approximately minimised by von Mises states, which play the role of squeezed states for quantum rotor and allow optimal detection of quantum states at the ultimate limits.
@@ I Introduction
Quantum mechanics imposes rules, which are establishing sophisticated network of interconnected and subtle conditions, which are distinguishing the realm of quantum effects from the classical world, and impose ultimate bounds on the precision of involved variables. This is, for instance, the case of the celebrated Heisenberg uncertainty relation for the canonical pair of position and momentum of a (quantum) particle. The uncertainty relation states that, loosely speaking, the momentum and position of a particle cannot be measured precisely at the same time. Interestingly, this concept generalizes to other pairs of canonical observables, for instance the quadrature operators of the electromagnetic field. In this case, the real part of the field can take the role of canonical position, whereas the imaginary part can be considered as the canonical momentum. These and similar examples have laid the foundations of quantum optics and quantum information processing. The important milestones on this route built systematically over one century were uncertainty relations, coherent (squeezed) states, Einstein-Podolsky-Rosen (EPR) states [1], Arthurs-Kelly concept of simultaneous detection of non-commuting variables [2, 3], concept of Bell-like measurement and representation of quantum state in phase-space pioneered by Wigner [4], Husimi [5] or Glauber [6]. On the top of those fundamental concepts there are several valuable protocols allowing process quantum information such as quantum teleportation [7] or quantum cryptography [8]. These examples can be cast into the same formal framework since they obey the commutation rule of the Lie algebra of the Heisenberg group, $[x,p]=i\hbar$, and observables are related by continuous Fourier transformation. Similar structure is associated with conservation of angular momentum, where operator of angular momentum $L$ and exponential angle operator $E=e^{-i\phi}$ obey the Lie algebra $\mathfrak{e}$(2) of the Euclidean group in the plane $E(2)$, $[E,L]=E$ [9, 10]. Such observables are related by the Discrete-Continuous Fourier transformation with the all consequences for quantum metrology and phase-space representation [11]. In spite of certain similarities between both the models, there are also significant differences stemming from the fact that formalism of quantum rotor represents the model beyond the harmonic oscillator [12].
@@
The purpose of this Article is driven by possible applications and has two objectives - to investigate and compare possible performance measures used for quantification of uncertainties of complementary variables, and to identify the physical models described by the operators comprising the $\mathfrak{e}$(2) algebra. The lack of saturable uncertainties is one of the reasons why the angular variable was considered as ”problematic”. For harmonic oscillator both the constructions adopting either variational approach or commutation rule tend to the same uncertainty relation. For the quantum rotor there is an option to use dispersion or the measure constructed from the covariance matrix of the exponential operator $E$. Though the uncertainties are not identical, they will yield consistent metrological consequences for optimal measurement. We provide in this paper a remarkable unified interpretation of both measures by extending the $\mathbb{R}^{2}$ plane, which is the obvious setting for the analysis of systems with one angle and one angular momentum, to the full $\mathbb{R}^{3}$ space. This interpretation comes by considering the measures as moments of inertia of an inhomogeneous ring about axes passing through its center of mass, but with rotations no longer necessarily limited to the original $\mathbb{R}^{2}$ plane. In this formulation, the moment of inertia tensor is given by a covariance matrix. Although this extension to 3D is apparently ad hoc, it is extremely advantageous from a metrological perspective as it can be used to formulate new tighter bounds if there is an experimental demand for this. We specify several physical models documenting the interdisciplinary role of quantum rotor: superconducting circuits with Josephson junction - a promising platform for quantum computation. Here the canonical pair of number of Cooper pairs tunneling through Josephson junction and magnetic flux in the super-conducting circuits [13, 14], shaping of optical pulses in 1D described by the Fourier series. The canonical pair corresponds to discrete modal index and continuous time variable. As the last example, let us mention beams with orbital angular momentum [15, 16, 17], which provide exquisite experimental platform for quantum information processing in higher dimension with several pivotal experiments [18, 19]. All these examples of quantum rotor-like systems may profit from theory which allows to quantify complementarity of conjugated variables of angular momentum and exponential angle variable.
@@
Historically the concept of angular momentum and angle is regarded as ”problematic” from several reasons, some of which we mention here. The misunderstanding can be related to periodicity, due to which the variance is not a good statistical measure since it is not shift invariant. Second, the uncertainties can be relatively easily formulated for sine or cosine of angular variable, as it was done in early days of quantum optics [20, 21]; however these inequalities cannot be saturated simultaneously and therefore extremal state cannot be fully exploited for metrological purposes. As a final example, the problem of angle and angular momentum was plagued by possible ambiguities in theoretical formulation; indeed, the quantum mechanics on the circle can be consistently formulated by means of variables where conjugated variables are combined in rather nontrivial ways [22]. Here the extremal states are the so called ”wrapped Gaussian states” - eigenstates of the operator $X=Ee^{-L-1/2}$, which however does not close on an algebra when combined with its conjugate operator. The thorough overview of theoretical concepts linked to angular momentum and angular variable can be found in [9, 23]. For all these reasons the quantum rotor was not considered in applications on an equal footing with the harmonic oscillator, though there are profound analogies in the mathematical description, stemming from the similarities between Fourier transformation and Fourier series [11].
@@
In Sec. II we review concepts for saturable uncertainty relations and extremal states for quantum rotor formulated in various contexts within several decades. Special attention will be payed to arguments based on variational principle and Robertson-like approach. In comparison with analogous theory for quadrature operators, there are subtle but substantial differences in case of quantum rotor. In Sec. III we unify both approaches on an equal footing introducing the moment of inertia. Importantly such a formulation not only unifies existing approaches but gives an opportunity to formulate new and tight uncertainties tailored to possible applications. In Sec. IV we extend the formulation to the problem of simultaneous optimal detection of non-commuting variables, which may find applications extended beyond its mechanical interpretation. Sec. V provides valuable examples where the generic theory developed here can be used including vortex beams, qubits in super-conducting circuits or optimal pulse shaping. Conclusions in Sec. VI summarize all the results stressing the metrological meaning of extremal states as fully fledged minimum uncertainty states for possible metrological applications. Technical calculations related to simplified derivation for generic moment of inertia and ultimate uncertainty relations for correlation matrix are reported in appendices.
@@ II Uncertainty relations for quantum rotor
    We can consider this as a Hamiltonian of the harmonic oscillator. Its minimum mean value is reached for state corresponding to the projection into the lowest eigenvalue state - squeezed vacuum state, displaced by all possible values $\langle x\rangle,\langle p\rangle.$ Such states provide over-complete set of (generalized) coherent states.
@@
    On the other hand the same solution can be obtained by an alternative way by adopting the uncertainties
@@
Here, we are interested in a pair of angular momentum operator $L=-i\partial_{\phi}$ and unitary exponential operator $E=e^{-i\phi}$ satisfying the commutation rule [9, 10]
@@
of Euclidean algebra $\mathfrak{e}$(2). In order to formulate uncertainties for angular momentum and angular variable $(L,\phi),$ notice that the standard variance $(\Delta\phi)^{2}$ is not a good uncertainty measure since it is not shift invariant. Statistical dispersion
@@
represents the simplest choice of the figure of merit including higher order moments of the ”angular variable”, not just its variance $(\Delta\phi)^{2}.$ However, as will be seen from the following the angular variable itself will be avoided in favour of exponential angle operator.
@@
The bound $B(D)$ is depicted by the solid red line in Fig. 3. However, the Mathieu ground state can be very closely approximated by the von Mises state $|m,\alpha,\kappa\rangle$ with $m=0$ and $\alpha=0$, where in $\phi$-representation one has [26, 24]
@@
$\kappa\geq 0$ being a parameter analogous to squeezing in close analogy with formulation for quadrature operators. Differences between ground Mathieu and von Mises states are so small (compare solid red line and dashed blue line in Fig. 3) that hardly can be observed by current technology and due to this, we will replace this analytically intractable bound $B(D)$ by its good approximation in terms of von Mises states. Explicit form can be easily found with the help of formulae derived in [11] in parametric form depending on $\kappa$
@@
There are physical reasons why Mathieu and von Mises ground states are so closely related. Whereas the Mathieu ground state minimises the uncertainty product for dispersion and variance of angular momentum, von Mises states are extremal states for uncertainty product of angular momentum and exponential angle operator. Indeed the commutation relations for rotated sine and cosine operators
@@
which is saturated by the von Mises states (7) - the solution of the operator equation
@@
$\langle(\Delta L)^{2}\rangle\geq\frac{1}{4}\underset{\|{\bf x}\|=1}{\mbox{max}}\frac{\left({\bf c}^{\rm T}{\bf x}\right)^{2}}{\left({\bf x}^{\rm T}{\bf\Gamma}{\bf x}\right)}=\frac{1}{4}\frac{\left({\bf c}^{\rm T}{\bf x}_{\rm o}\right)^{2}}{{\bf x}_{\rm o}^{\rm T}{\bf\Gamma}{\bf x}_{\rm o}}=\frac{1}{4}{\bf c}^{\rm T}{\bf\Gamma}^{-1}{\bf c},$ (16)
@@
$\langle(\Delta L)^{2}\rangle\stackrel{{\scriptstyle 1}}{{\geq}}\frac{1}{4}\frac{\left({\bf c}^{\rm T}{\bf x}_{\rm o}\right)^{2}}{\left({\bf x}_{\rm o}^{\rm T}{\bf\Gamma}{\bf x}_{\rm o}\right)}\stackrel{{\scriptstyle 2}}{{\geq}}\frac{1}{4}\frac{\|{\bf c}\|^{2}}{{\bf m}^{\rm T}{\bf\Gamma}{\bf m}}\stackrel{{\scriptstyle 3}}{{\geq}}\frac{1}{4}\frac{\|{\bf c}\|^{2}}{\gamma_{+}}\stackrel{{\scriptstyle 4}}{{\geq}}\frac{1}{4}\frac{\|{\bf c}\|^{2}}{D^{2}},$ (17)
@@
The denominators ${\bf x}_{\rm o}^{\rm T}{\bf\Gamma}{\bf x}_{\rm o}$, ${\bf m}^{\rm T}{\bf\Gamma}{\bf m}$ and $\gamma_{+}$ appearing on the right-hand sides of the chain of inequalities (17) represent possible alternative uncertainty measures of angular variable constructed from elements of the covariance matrix (13). These measures coincide for von Mises states, ${\bf x}_{\rm o}^{\rm T}{\bf\Gamma}{\bf x}_{\rm o}={\bf m}^{\rm T}{\bf\Gamma}{\bf m}=\gamma_{+}$, and all the inequalities $1,2$ and $3$ reduce to equalities. However, the measures are not the same for other than extremal states, since they are forged from different parameters of covariance matrix. The least one is ${\bf x}_{\rm o}^{\rm T}{\bf\Gamma}{\bf x}_{\rm o}$ and it is on the RHS of inequality $1$ but the simplest one is given on the RHS of inequality $2$ and it was heuristically derived in [11], ${\bf m}^{\rm T}{\bf\Gamma}{\bf m}=\langle S_{-\mbox{arg}\langle E\rangle}^{2}\rangle$. Note finally that the quantities ${\bf x}_{\rm o}^{\rm T}{\bf\Gamma}{\bf x}_{\rm o}$ and ${\bf m}^{\rm T}{\bf\Gamma}{\bf m}$ are ill-defined for states with $\langle E\rangle=0$, i.e., ${\bf c}=(0,0)^{\rm T}$, in which case we define both of them as $\gamma_{-}=(1-|\langle E^{2}\rangle|)/2$.
In the next section we show that concept of moment of inertia unifies all the concepts analysed above, and allows to design tighter uncertainty relations.
III Moments of inertia as the angular uncertainty measures

### 13 Mar 2024 v2

@@ Abstract
Quantum rotor represents the second simplest model after the harmonic oscillator with profound interdisciplinary but yet unexplored implications. It overreaches its mechanical meaning with significant consequences and promising applications in, e.g., singular optics, super-conductive circuits with Josephson junction or optimal pulse shaping in time-frequency domains. Quantification of complementarity between angular momentum and angular variable is essential for quantum metrology. Whereas the natural measures for position and momentum are variances, the uncertainty associated with angular variable can be linked to moments of inertia of the unit ring about axes passing through its center of mass. This interpretation unifies variants for saturable uncertainty relations and explains ambiguities in choice of possible uncertainty measures associated with sine and cosine functions of angular variable. Special attention is payed to uncertainty products which are exactly or approximately minimised by von Mises states and play the role of squeezed states for quantum rotor allowing optimal detection of quantum states at the ultimate limits.
@@ I Introduction
Quantum mechanics imposes rules, which are establishing sophisticated network of interconnected and subtle conditions, which are distinguishing the realm of quantum effects from the classical world and provides ultimate bounds on the precision of involved variables. This is, for instance, the case of the celebrated Heisenberg uncertainty relation for the canonical pair of position and momentum of a (quantum) particle. The uncertainty relation states that, loosely speaking, the momentum and position of a particle cannot be measured precisely at the same time. Interestingly, this concept generalizes to other pairs of canonical observables, for instance the quadrature operators of the electromagnetic field. In this case, the real part of the field can take the role of canonical position, whereas the imaginary part can be considered as the canonical momentum. These and similar examples have laid the foundations of quantum optics and quantum information processing. The important milestones on this route built systematically over one century were uncertainty relations, coherent (squeezed) states, Einstein-Podolsky-Rosen (EPR) states [1], Arthurs-Kelly concept of simultaneous detection of non-commuting variables [2, 3], concept of Bell-like measurement and representation of quantum state in phase-space pioneered by Wigner [4], Husimi [5] or Glauber [6]. On the top of those fundamental concepts there are several valuable protocols allowing process quantum information such as quantum teleportation [7] or quantum cryptography [8]. These examples can be cast into the same formal framework since they obey the commutation rule of the Lie algebra of the Heisenberg group, $[x,p]=i\hbar$, and observables are related by continuous Fourier transformation. Similar structure is associated with conservation of angular momentum, where operator of angular momentum $L$ and exponential angle operator $E=e^{-i\phi}$ obey the Lie algebra $\mathfrak{e}$(2) of the Euclidean group in the plane $E(2)$, $[E,L]=E$ [9, 10]. Such observables are related by the Discrete-Continuous Fourier transformation with the all consequences for quantum metrology and phase-space representation [11]. In spite of certain similarities between both the models, there are also significant differences stemming from the fact that formalism of quantum rotor represents the model beyond the harmonic oscillator [12].
@@
The purpose of this Article is driven by possible applications and has two objectives - to investigate and compare possible performance measures used for quantification of uncertainties of complementary variables, and to identify the physical models described by the operators comprising the $\mathfrak{e}$(2) algebra. The lack of saturable uncertainties is one of the reasons why the angular variable was considered as problematic. For harmonic oscillator both the constructions adopting either variational approach or commutation rule tend to the same uncertainty relation. For the quantum rotor there is an option to use dispersion or the measure constructed from the covariance matrix of the exponential operator $E$. Though the uncertainties are not identical, they will yield consistent metrological consequences for optimal measurement. In this paper we provide a remarkable unified interpretation of both measures by extending the $\mathbb{R}^{2}$ plane, which is the obvious setting for the analysis of systems with one angle and one angular momentum, to the full $\mathbb{R}^{3}$ space. This interpretation comes by considering the measures as moments of inertia of an inhomogeneous ring about axes passing through its center of mass, but with rotations no longer necessarily limited to the original $\mathbb{R}^{2}$ plane. The moment of inertia tensor is given by a covariance matrix. Although this extension to 3D is apparently ad hoc, it is extremely advantageous from a metrological perspective as it can be used to formulate new tighter bounds if there is an experimental demand for this. Besides numerous mechanical implementations of quantum rotors [13, 14], we specify several physical models documenting the interdisciplinary role of quantum rotor. Superconducting circuits with Josephson junction represent a promising platform for quantum computation. Here the canonical pair is given by number of tunneling Cooper pairs and the phase over Josephson junction [15, 16]. For shaping of optical pulses in 1D described by the Fourier series the canonical pair corresponds to discrete modal index and continuous time variable. As the last example, beams with orbital angular momentum [17, 18, 19] provide exquisite experimental platform for quantum information processing in higher dimension with several pivotal experiments [20, 21]. All these examples of quantum rotor-like systems may profit from theory which allows to quantify complementarity of conjugated variables of angular momentum and exponential angle variable.
@@
Historically the concept of angular momentum and angle was regarded as unsettled from several reasons, some of which we mention here. The misunderstanding can be related to periodicity, due to which the variance is not a good statistical measure since it is not shift invariant. Second, the uncertainties can be relatively easily formulated for sine or cosine of angular variable, as it was done in early days of quantum optics [22, 23]; however these inequalities cannot be saturated simultaneously and therefore such extremal states cannot be fully exploited for metrological purposes. As a final remark, the problem of angle and angular momentum was plagued by possible ambiguities in theoretical formulation; indeed, the quantum mechanics on the circle can be consistently formulated by means of variables where conjugated variables are combined in rather nontrivial ways [24]. Here the extremal states are the so called ”wrapped Gaussian states” - eigenstates of the operator $X=Ee^{-L-1/2}$, which however does not close on an algebra when combined with its conjugate operator. The thorough overview of theoretical concepts linked to angular momentum and angular variable can be found in [9, 25]. For all these reasons the quantum rotor was not considered in applications on an equal footing with the harmonic oscillator, though there are profound analogies in the mathematical description, stemming from the similarities between Fourier transformation and Fourier series [11].
@@
In Sec. II we review concepts for saturable uncertainty relations and extremal states for quantum rotor formulated in various contexts within several decades. Special attention will be payed to arguments based on variational principle and Robertson-like approach. In comparison with analogous theory for quadrature operators, there are subtle but substantial differences in case of quantum rotor. In Sec. III we unify both approaches on an equal footing introducing the moment of inertia. Importantly such a formulation not only unifies existing approaches but gives an opportunity to formulate new and tight uncertainties tailored to possible applications. In Sec. IV we extend the formulation to the problem of simultaneous optimal detection of non-commuting variables, which may find applications extended beyond its mechanical interpretation. Sec. V provides valuable examples where the generic theory developed here can be used including vortex beams, qubits in super-conducting circuits or optimal pulse shaping. Conclusions in Sec. VI summarize all the results stressing the metrological meaning of extremal states as fully fledged minimum uncertainty states for possible metrological applications. Technical calculations related to simplified derivation for generic moment of inertia, ultimate uncertainty relations for correlation matrix, and optimal simultaneous measurement are reported in appendices.
@@ II Uncertainty relations for quantum rotor
    We can consider this as a Hamiltonian of the harmonic oscillator. Its minimum mean value is reached for a state corresponding to the projection into the lowest eigenvalue state - squeezed vacuum state, displaced by all possible values $\langle x\rangle,\langle p\rangle.$ Such states provide over-complete set of (generalized) coherent states.
@@
    On the other hand the same solution can be obtained by an alternative way using the inequalities
@@
Quantum rotor is characterized by angular momentum operator $L=-i\partial_{\phi}$ and unitary exponential operator $E=e^{-i\phi}$ satisfying the commutation rule [9, 10]
@@
of Euclidean algebra $\mathfrak{e}$(2). In order to formulate uncertainties for angular momentum and angular variable $(L,\phi),$ notice that the standard variance $\langle(\Delta\phi)^{2}\rangle$ is not a good uncertainty measure since it is not shift invariant. Statistical dispersion
@@
represents the simplest choice of the figure of merit including higher order moments of the ”angular variable”, not just its variance $\langle(\Delta\phi)^{2}\rangle$. In the following the angular variable itself will be avoided in favour of exponential angle operator.
@@
The bound $B(D)$ is depicted by the solid red line in Fig. 3. Importantly the Mathieu ground state can be very closely approximated by the von Mises state $|m,\alpha,\kappa\rangle$ with $m=0$ and $\alpha=0$, where in $\phi$-representation one has [28, 26]
@@
$\kappa\geq 0$ being a parameter resembling squeezing in close analogy with formulation for quadrature operators. Differences between ground Mathieu and von Mises states are so small (compare solid red line and dashed blue line in Fig. 3) that hardly can be observed by current technology and due to this, we will replace this analytically intractable bound $B(D)$ by its good approximation in terms of von Mises states. Explicit form can be easily found with the help of formulae derived in [11] in parametric form depending on $\kappa$
@@
There are physical reasons why Mathieu and von Mises ground states are so closely related. Whereas the Mathieu ground state minimises the uncertainty product for dispersion and variance of angular momentum, von Mises states are extremal states for uncertainty product of angular momentum and exponential angle operator. Indeed the commutation relation for rotated sine and cosine operators
@@
which is saturated by the von Mises states (7) as the solution of the operator equation
@@
$\langle(\Delta L)^{2}\rangle\geq\frac{1}{4}\underset{\|{\bf x}\|=1}{\mbox{max}}\frac{\left({\bf c}^{\rm T}{\bf x}\right)^{2}}{{\bf x}^{\rm T}{\bf\Gamma}{\bf x}}=\frac{1}{4}\frac{\left({\bf c}^{\rm T}{\bf x}_{\rm o}\right)^{2}}{{\bf x}_{\rm o}^{\rm T}{\bf\Gamma}{\bf x}_{\rm o}}=\frac{1}{4}{\bf c}^{\rm T}{\bf\Gamma}^{-1}{\bf c},$ (16)
@@
$\langle(\Delta L)^{2}\rangle\stackrel{{\scriptstyle 1}}{{\geq}}\frac{1}{4}\frac{\left({\bf c}^{\rm T}{\bf x}_{\rm o}\right)^{2}}{{\bf x}_{\rm o}^{\rm T}{\bf\Gamma}{\bf x}_{\rm o}}\stackrel{{\scriptstyle 2}}{{\geq}}\frac{1}{4}\frac{\|{\bf c}\|^{2}}{{\bf m}^{\rm T}{\bf\Gamma}{\bf m}}\stackrel{{\scriptstyle 3}}{{\geq}}\frac{1}{4}\frac{\|{\bf c}\|^{2}}{\gamma_{+}}\stackrel{{\scriptstyle 4}}{{\geq}}\frac{1}{4}\frac{\|{\bf c}\|^{2}}{D^{2}},$ (17)
@@
The denominators ${\bf x}_{\rm o}^{\rm T}{\bf\Gamma}{\bf x}_{\rm o}$, ${\bf m}^{\rm T}{\bf\Gamma}{\bf m}$ and $\gamma_{+}$ appearing on the right-hand sides of the chain of inequalities (17) represent possible alternative uncertainty measures of angular variable constructed from elements of the covariance matrix (13). These measures coincide for von Mises states, ${\bf x}_{\rm o}^{\rm T}{\bf\Gamma}{\bf x}_{\rm o}={\bf m}^{\rm T}{\bf\Gamma}{\bf m}=\gamma_{+}$, and all the inequalities $1,2$ and $3$ reduce to equalities. However, the measures are not the same for other than extremal states, since they are forged from different parameters of covariance matrix. The least one is ${\bf x}_{\rm o}^{\rm T}{\bf\Gamma}{\bf x}_{\rm o}$ and it is on the RHS of inequality $1$ but the simplest one is given on the RHS of inequality $2$ and it was heuristically derived in [11], ${\bf m}^{\rm T}{\bf\Gamma}{\bf m}=\langle S_{-\mbox{arg}\langle E\rangle}^{2}\rangle$. Note finally that the quantities ${\bf x}_{\rm o}^{\rm T}{\bf\Gamma}{\bf x}_{\rm o}$ and ${\bf m}^{\rm T}{\bf\Gamma}{\bf m}$ are ill-defined for states with $\langle E\rangle=0$, i.e., ${\bf c}=(0,0)^{\rm T}$, in which case we define both of them as $\gamma_{-}=(1-|\langle E^{2}\rangle|)/2$. In the next section we show that concept of moment of inertia unifies all the uncertainties analysed above, and allows to design even tighter uncertainty relations.
III Moment of inertia as the angular uncertainty measure